OpenStudy (anonymous):

10.) If a triangle has sides of lengths 6 sqrt(2), 6 sqrt(2), and 12, is it a right triangle? Yes No Sometimes Not enough information

OpenStudy (anonymous):

yes I think

OpenStudy (dumbcow):

we already told you how to find out??

OpenStudy (anonymous):

actually it is no

OpenStudy (dumbcow):

6sqrt(2) ^2 = 36*2=72 72+72=144 = 12^2 the answer is yes